OpenStudy (anonymous):

Rational equation problem: Student A can finish a job in 6 hours while it would take 8 hours for Student B to do the same job. If they work together, how long would it take them to finish the job?

jhonyy9 (jhonyy9):

student A ---- 6 hours student B ---- 8 hours ------------------- together ----- (6+8)/2 = 14/2 = 7 hours